# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 4.1 Explainable Artificial Intelligence (XAI)
A formal definition of XAI is given in section 1.2. This section explores in more detail of why, what, and how of XAI. The omnipresence of AI dictates how powerful it is and has become a part of our lives. It is being applied in making critical decisions, especially in the field of health (jimma2023artificial, ; afnan2021interpretable, ) and forensic sciences (galante2023applications, ). There are certain issues facing AI when applied in these domains. One of the major issues is trust. For example, when a forensic analyst identifies the suspect of a child exploitation case with the help of the reports generated by an AI system. At the time this suspect is presented in a court of law with the evidence, will the court be able to give a judgment based on these findings? To trust these findings, the court will require an explanation of these findings, and the AI methods used to reach the findings will need to be interpreted and explained. This is one of the core requirements of any forensic analysis during a cyber crime investigation to be accepted and presented in a court of law. The main purpose of XAI is to develop trust in AI and play a central role in a successful legal CF analysis.
There are three basic approaches to generating explanations: (1) Using simple models, such as decision trees and other rule-based approaches. The output of these models depends on a small set of rules that are easy to explain to the system’s user. (2) Feature selection and analysis, focusing on a small set of features. This makes it easy to understand the relationship between input and output even if the model is complex. (3) The most prevalent approach uses external mechanisms to help describe the inner workings of a black-box AI system. One of the methods in this approach creates a proxy model, such as a decision tree, to understand and then explain the original model. One other method perturbs the model by querying the model with different inputs to assess which inputs are important. Popular XAI algorithms within this approach are LIME (Local Interpretable Model-Agnostic Explanations) (ribeiro2016should, ) and SHAP (SHapley Additive exPlanation) (lundberg2020local, ; lundberg2017unified, ). National Institute of Standards and Technology introduced four principles (phillips2021four, ) and proposed that all XAI systems should adhere to these principles.
@@
    Explanation – A system is explainable if its outcome can be reasoned logically, intelligently, and methodically. Every AI system in practice has a different way of explaining itself. Therefore, to keep it generic this principle does not put any condition of correctness or impose any quality metrics. There are two types of XAI models, self-explainable and post-hoc explainable. Self-Explainable as the name suggests explain themselves. The model can be explained globally as well as locally. Each decision on an input can be explained, also called local explanation, and an explanation can also be provided for all the inputs as a whole, also called global explanation. Most of the common self-explainable models are decision trees and regression. These are some of the simplest AI models and are designed to be more meaningful to humans. Post-Hoc Explainable as the name suggests explains after the output is seen. These explanations are rendered by other software tools and can be used on the algorithms without knowing their inner workings, provided the outputs can be queried over the inputs. These explanations produce feature importance scores, rule sets, heatmaps, or natural language representations. This helps a user quantify the importance of features and explain the decisions produced by the model. Some researchers (caruana2020intelligible, ; dovsilovic2018explainable, ; linardatos2020explainable, ) argue the trade-off of simplicity with accuracy and state that a self-explainable model is less accurate than a post-hoc model. There are other researchers (rudin2019stop, ; rudin2019we, ; afnan2021interpretable, ) that disagree with this statement.
@@
    Meaningful – An AI system is meaningful if its explanation is understandable. Stakeholders of the AI system play an important role in this principle. To consider an explanation meaningful depends on the stakeholders’ knowledge, experiences, relationship to the system, and various other psychological factors. The other major factor is the purpose of the explanation, such as it will be different for legal forensics than for a business purpose. This makes the explanation intelligible to the intended audience. There are some challenges in making the explanation meaningful for a variety of different audiences. For example, a forensic scientist explains the evidence to a layperson, e.g., a member of a jury. Such an explanation in general is difficult to understand and may mislead a layman reader, and therefore do not fulfill the criteria for being meaningful (jackson2015communicating, ). Some of the challenges (jackson2015communicating, ; lombrozo2006structure, ; mueller2019explanation, ) for producing a meaningful explanation are individual differences of the users about a meaningful explanation, context of an explanation, and expectation of users about different explanation types. An XAI system should provide meaningful information appropriate to the context and understandable by the intended users.
@@
    Explanation Accuracy – This is the ability of the explanation to accurately describe how the AI system reaches its conclusion. This is different than the accuracy of an AI algorithm and system. The level of detail in the explanation is also important. For some audiences simple explanations suffice, others may require more details to better understand the outcome. An inventor may explain her/his invention with substantial technical details to colleagues. The same explanation will likely be presented differently to untrained friends and parents. The accuracy of explanations depends on how meaningful they are to certain audiences. Therefore, meaningfulness and explanation accuracy are both associated. There are metrics for measuring the accuracy of an AI system (japkowicz2011evaluating, ; phillips2000feret, ; phillips2007frvt, ; phillips2008iris, ; przybocki2007nist, ; reynolds20172016, ; hossin2015review, ), such as receiver operating characteristic curve and area under the curve (Introduction-ROC-Analysis, ), precision-recall curve (Precision-Recall, ), and cost curves (Cost-Curves, ) etc. Besides these, there is also work (hoffman2018metrics, ; rosenfeld2021better, ; sovrano2023objective, ; lopes2022xai, ) in progress on developing metrics for evaluating the performance of an XAI system. How to measure the accuracy of an XAI system? There are certain criteria, such as, how good (precise and clear) is the explanation, how satisfied it is for the respective audience, and how trustful (justifiable) is the explanation. A detailed taxonomy of the XAI system’s evaluation methods is given in (lopes2022xai, ).
@@
    Knowledge Limits – This principle makes sure that the AI system operates within the limits of the system’s knowledge and design. Any operation or outcome that is out of these limits produces a false judgment and hence is not trusted. This helps the system to dismiss misleading, and ambiguous outputs/judgments and hence increases the trust in an AI system. There are two ways an AI system can exceed its knowledge limits. One is if the query or operation is outside the domain of the AI system. As an example, in CF successful file fragment classification becomes challenging for an AI algorithm when an image is embedded in a file that itself is not an image, e.g., an image embedded in a PDF file (alam2023sift, ). In this case, if the AI algorithm classifies the fragment as an image rather than a PDF file without any explanation then the system using the AI algorithm will not be able to dismiss this as a misleading output. But if the AI algorithm knows its limitations and explains then the AI system will be able to dismiss this misleading output. The second is if the confidence of the output is too low. Revisiting the above example, classifying a fragment in a multi-class (several file types) is also a challenging problem (alam2023sift, ). In this case, if the AI algorithm classifies a fragment into more than one class and randomly chooses one class out of these as an output, then this will be a misleading output. An AI system with the knowledge of these limitations will not trust such output.
@@ 4.1.1 Classification of XAI Methods
We divide and classify the XAI methods by the scope of their explanation as shown in Figure 4. The scope of an explanation refers to its range and extent. It can be either local or global. Both local and global can be intrinsic and post-hoc. By definition, intrinsic methods are model-specific, and post-hoc methods are generally model-agnostic. This section presents some of the popular local and global methods. For a comprehensive coverage on this specific subject readers are referred to (vilone2020explainable, ; guidotti2018survey, ; arrieta2020explainable, ).
@@
    Local Explanations – The scope of these explanations is a single input or a subset of inputs. The most common is the explanation for the output of a single input point. Some of the local explanation algorithms are LIME (ribeiro2016should, ), SHAP (lundberg2020local, ; lundberg2017unified, ), Saliency Map (iyer2018transparency, ; selvaraju2017grad, ), and Counterfactual (wachter2017counterfactual, ). LIME explains each prediction of the original model by building a proxy self-interpretable model, such as linear models or decision trees. This new model is trained on a new dataset containing the interpretable representations of the original data. The original model can be explained by the weights given to features in the new model. For example, a model predicts that a patient has flu. LIME in the explanation highlights the symptoms with relative weights, such as sneezing and headache are given more weight than no fatigue. A doctor can easily decide with the help of these explanations. SHAP also builds and uses a simpler model that is the interpretable approximation of the original model. It provides feature importance (contribution of each feature to the prediction result) by computing the Shapley value for each feature based on game theory. Saliency Map also known as sensitivity map or heat map explains each prediction by computing an attribute value for each feature based on its degree of influence on the prediction result. A saliency map is a visual explanation of a model where contributions can be visualized as heat maps. Counterfactual explanations indicate what amount of change in the input can modify the output. The objective is to minimize the new input value and bring it as close to the original input as possible. The goal is to maximize this objective function by a local search. This function depends on the endpoint loss, i.e., how close its value is to the desired class and the distance between the origin and the new value. Counterfactual explanations are similar to adversarial examples but they differ in purpose. Unlike adversarial examples where the purpose is to evade a classifier, they facilitate understanding of a model by providing explanations.
@@
    Global Explanations – The scope of these explanations is the entire algorithm. Some of the global explanation algorithms are PDP (Partial Dependence Plot) (friedman2001greedy, ), ICE (Individual Conditional Expectation) (goldstein2015peeking, ), GSA (Global Sensitivity Analysis) (cortez2011opening, ; cortez2013using, ), SP-LIME (Submodular Pick LIME) (ribeiro2016should, ). PDP explanations indicate the marginal change of the output when there is a change in the feature value. The PDP focuses on the overall average instance rather than a specific one. It reveals the relationship between a feature and the prediction results. ICE plots are the graphs with line charts that represent the functional relationship between a feature and the predicted response. All the other features are kept fixed, only changing the value of the feature under analysis. GSA method perturbs the input features through their range of values and quantifies its effects on the predictions of a given model. SP-LIME, a variant of LIME, chooses the most relevant local LIME explanations to summarize and provide a global explanation.
@@ 4.2 Last Ten Years of XAI
To make readers familiar with the study carried out in the paper, we present a critical and short review of the last ten years of work carried out in XAI. We exclude papers that explore the application of XAI in CF because we discuss in detail such applications in section 6. A detailed survey on this topic is out of the scope of this paper. For a detailed survey, interested readers are referred to (adadi2018peeking, ; ali2023explainable, ).
Wang et al. (wang2015falling, ) introduced a Bayesian framework for learning Falling Rule Lists (FRL), a type of interpretable predictive model designed for decision-making in healthcare applications. FRLs consist of an ordered list of if-then rules, prioritizing high-risk patients for classification first. The Bayesian approach enhances interpretability, providing a transparent decision-making process, which could prove beneficial in medical contexts, aligning with the growing importance of trust and understanding in high-stakes decisions. Su et al. (su2015interpretable, ) employed optimization-based algorithms for learning two-level Boolean rules, focusing on achieving a balance between classification accuracy and rule simplicity. The introduced algorithms, considering both classification accuracy and human interpretability, demonstrate superior performance, contributing to interpretable models with improved performance in AI contexts. Letham et al. (letham2015interpretable, ) introduced Bayesian Rule Lists (BRL) a generative model for decision lists, aiming to strike a balance between predictive accuracy and interpretability. The method, driven by a hierarchical prior over permutations of rules, produces decision lists with high posterior probability, offering both accuracy and interpretability. This showcases the potential for interpretable models in personalized medicine, emphasizing the importance of concise and trustworthy rules for effective communication with domain experts. Hara et al. (hara2016making, ) proposes a post-processing method to enhance the interpretability of Additive Tree Models (ATMs) (cui2015optimal, ) by approximating them with simpler, more human-interpretable models. The challenge addressed is the high number of regions that ATMs typically create, hindering interpretability.
@@
Mishra et al. (mishra2017local, ) extend the LIME technique to music content analysis (MCA) named SLIME, introducing three versions of explanations based on temporal, frequency, and time-frequency segmentation. SLIME provides insights into model behavior, revealing limitations in the decision tree model’s generalization despite high accuracy and demonstrating model-agnostic explanations for the neural network aligning with model-dependent saliency maps in many cases. Fong et al. (fong2017interpretable, ) propose a model-agnostic framework for learning explanations for any black box algorithm, with a focus on identifying the image regions most responsible for classifier decisions. The proposed method employs explicit and interpretable image perturbations, addressing the limitations of heuristic-based image saliency methods. The framework contributes to interpretable and testable explanations, revealing insights into neural network fragility and susceptibility to artifacts. Ribeiro et al. (ribeiro2018anchors, ) propose a model-agnostic explanation system called Anchors, which provides high-precision rules in the form of if-then statements to elucidate the behavior of complex models. Anchors efficiently compute explanations for any black-box model with high probability, showcasing their flexibility across various domains and tasks. A user study demonstrates that Anchors enable users to predict model behavior on unseen instances with higher precision and less effort compared to existing model-agnostic explanation techniques. Guidotti et al. (guidotti2018local, ) introduce LORE (for LOcal Rule-based Explanations), a model-agnostic method for explaining black box decisions, focusing on providing interpretable and faithful explanations for specific instances. LORE employs a genetic algorithm to generate a synthetic neighborhood and then builds a local interpretable predictor in the form of a decision tree. The resulting explanation includes a decision rule and a set of counterfactual rules. LORE outperforms existing methods in terms of explanation quality and accuracy in mimicking the black box, addressing the ethical issue of lacking transparency in decision systems.
@@
Joo et al. (joo2019visualization, ) investigate the application of Grad-CAM to Deep Reinforcement Learning (DRL) in the context of Atari games. By revealing the significance of the CNN layer, the research enhances our understanding of how AI agents make decisions while playing Atari games through the application of Grad-CAM to DRL. Gramegna et al. (gramegna2020buy, ) propose an XAI model for deciphering customer decisions in the non-life insurance domain. The authors integrate XGBoost with Shapley Values that enhances behavioral segmentation, showcasing the potential of explainable machine learning to refine our understanding of customer behavior in the insurance industry. Cavaliere et al. (cavaliere2020parkinson, ) address the challenge of poor explainability in complex machine learning models, particularly in medical applications. Their research focused on automatic Parkinson’s disease detection, utilizing Grammar Evolution (GE) with Automatically Defined Functions. Parsa et al. (parsa2020toward, ) apply the eXtreme Gradient Boosting (XGBoost) machine learning technique to enhance traffic safety through rapid accident detection. The study demonstrated XGBoost’s effectiveness with high accuracy and detection rates, emphasizing the significance of traffic-related features, especially speed variations, through SHAP analysis. Feichtner et al. (feichtner2020understanding, ) aims to identify inconsistencies between developer-defined application behavior and permission usage in Android applications. It utilizes a deep neural network trained on descriptive texts to predict permission scores, with LIME employed for explainability analysis. De et al. (de2020explainable, ) propose a method to enhance the interpretability of decisions made by Deep Neural Networks (DNNs) by combining Cluster-TREPAN, a decision tree generation algorithm, with hidden-layer-clustering. The focus is on explaining predictions from neural networks, emphasizing the importance of interpretability and reliability. The proposed Cluster-TREPAN method outperforms LIME in performance, effectively capturing and explaining the information flow within DNNs, addressing a critical aspect for human acceptance of deep learning models.
@@
Szczepański et al. (szczepanski2021new, ) focused on enhancing the explainability of BERT-based fake news detectors, addressing the crucial need for understanding model decisions in real-life scenarios. Utilizing surrogate-type explainability methods, specifically LIME and Anchors, the study demonstrates the feasibility of enhancing interpretability in linguistic-based fake news detection models. Gite1 et al. (gite2021explainable, ) utilize an advanced combination of efficient machine learning techniques, specifically Long Short Term Memory (LSTM) for stock price prediction. The paper introduces an XAI element with LIME for transparent and meaningful insights into the model’s decision-making process. Naeem et al. (naeem2022explainable, ) address the critical issue of automated malware detection in IoT devices, leveraging deep learning techniques, and using Grad-CAM to understand and improve the overall performance. Sutton et al. (sutton2022artificial, ) Sutton et al. apply the Grad-CAM technique to enhance the diagnostic precision of ulcerative colitis, utilizing 8000 endoscopic images. While the choice of Grad-CAM provides visual insights into model learning, a more detailed discussion of the AI techniques used and the specific XAI methods employed would provide a comprehensive evaluation of the study’s contributions. Debjit et al. (debjit2022improved, ) work present an advanced machine learning framework for early COVID-19 detection, leveraging the Harris Hawks Optimization algorithm to optimize ML classifiers’ hyperparameters and SHAP values for model interpretation. Utilizing SHAP, the research offers transparency by elucidating feature importance, contributing to critical clinical insights in the domain of healthcare monitoring for COVID-19. Nordin et al. (nordin2023explainable, ) propose an explainable predictive model for suicide attempts, combining complex ensemble learning models, such as Random Forest and Gradient Boosting, with the SHAP method. This approach aims to enhance reliability in predicting suicide attempts while offering insights into critical factors for clinical decision-making. In the domain of mental health, the study highlights the superiority of Gradient Boosting with SHAP over Random Forest in terms of accuracy, underscoring the value of explainable machine learning techniques in healthcare.

### v3

@@ 4.1 Explainable Artificial Intelligence (XAI)
A formal definition of XAI is given in section 1.2. This section explores in more detail of why, what, and how of XAI. The omnipresence of AI dictates how powerful it is and has become a part of our lives. It is being applied in making critical decisions, especially in the field of health [41, 42] and forensic sciences [43]. There are certain issues facing AI when applied in these domains. One of the major issues is trust. For example, when a forensic analyst identifies the suspect of a child exploitation case with the help of the reports generated by an AI system. At the time this suspect is presented in a court of law with the evidence, will the court be able to give a judgment based on these findings? To trust these findings, the court will require an explanation of these findings, and the AI methods used to reach the findings will need to be interpreted and explained. This is one of the core requirements of any forensic analysis during a cybercrime investigation to be accepted and presented in a court of law. The main purpose of XAI is to develop trust in AI and play a central role in a successful legal CF analysis.
There are three basic approaches to generating explanations: (1) Using simple models, such as decision trees and other rule-based approaches. The output of these models depends on a small set of rules that are easy to explain to the system’s user. (2) Feature selection and analysis, focusing on a small set of features. This makes it easy to understand the relationship between input and output, even if the model is complex. (3) The most prevalent approach uses external mechanisms to help describe the inner workings of a black-box AI system. One of the methods in this approach creates a proxy model, such as a decision tree, to understand and then explain the original model. One other method perturbs the model by querying the model with different inputs to assess which inputs are important. Popular XAI algorithms within this approach are LIME (Local Interpretable Model-Agnostic Explanations) [44] and SHAP (SHapley Additive exPlanation) [45, 46]. The National Institute of Standards and Technology introduced four principles [47] and proposed that all XAI systems should adhere to these principles.
@@
    Explanation – A system is explainable if its outcome can be reasoned logically, intelligently, and methodically. Every AI system in practice has a different way of explaining itself. Therefore, to keep it generic, this principle does not put any condition of correctness or impose any quality metrics. There are two types of XAI models, self-explainable and post-hoc explainable. Self-Explainable as the name suggests, explain themselves. The model can be explained globally as well as locally. Each decision on an input can be explained, also called local explanation, and an explanation can also be provided for all the inputs as a whole, also called a global explanation. Most of the common self-explanatory models are decision trees and regression. These are some of the simplest AI models and are designed to be more meaningful to humans. Post-Hoc Explainable as the name suggests, explains after the output is seen. These explanations are rendered by other software tools and can be used on the algorithms without knowing their inner workings, provided the outputs can be queried over the inputs. These explanations produce feature importance scores, rule sets, heatmaps, or natural language representations. This helps a user quantify the importance of features and explain the decisions produced by the model. Some researchers [48, 49, 50] argue the trade-off of simplicity with accuracy and state that a self-explainable model is less accurate than a post-hoc model. There are other researchers [51, 52, 42] that disagree with this statement.
@@
    Meaningful – An AI system is meaningful if its explanation is understandable. Stakeholders of the AI system play an important role in this principle. Whether an explanation is meaningful depends on the stakeholders’ knowledge, experiences, relationship to the system, and various other psychological factors. The other major factor is the purpose of the explanation, such as it will be different for legal forensics than for a business purpose. This makes the explanation intelligible to the intended audience. There are some challenges in making the explanation meaningful for a variety of different audiences. For example, a forensic scientist explains the evidence to a layperson, e.g., a member of a jury. Such an explanation in general is difficult to understand and may mislead a layman reader, and therefore does not fulfill the criteria for being meaningful [53]. Some of the challenges [53, 54, 55] for producing a meaningful explanation are individual differences of the users about a meaningful explanation, the context of an explanation, and the expectation of users about different explanation types. An XAI system should provide meaningful information appropriate to the context and understandable by the intended users.
@@
    Explanation Accuracy – This is the ability of the explanation to accurately describe how the AI system reaches its conclusion. This is different than the accuracy of an AI algorithm and system. The level of detail in the explanation is also important. For some audiences, simple explanations suffice; others may require more details to better understand the outcome. An inventor may explain her/his invention with substantial technical details to colleagues. The same explanation will likely be presented differently to untrained friends and parents. The accuracy of explanations depends on how meaningful they are to certain audiences. Therefore, meaningfulness and explanation accuracy are both associated. There are metrics for measuring the accuracy of an AI system [56, 57, 58, 59, 60, 61, 62], such as receiver operating characteristic curve and area under the curve [63], precision-recall curve [64], and cost curves [65] etc. Besides these, there is also work [66, 67, 68, 69] in progress on developing metrics for evaluating the performance of an XAI system. How to measure the accuracy of an XAI system? There are certain criteria, such as, how good (precise and clear) the explanation is, how satisfied it is for the respective audience, and how trustworthy (justifiable) the explanation is. A detailed taxonomy of the XAI system’s evaluation methods is given in [69].
@@
    Knowledge Limits – This principle makes sure that the AI system operates within the limits of the system’s knowledge and design. Any operation or outcome that is out of these limits produces a false judgment and hence is not trusted. This helps the system to dismiss misleading and ambiguous outputs/judgments and hence increases the trust in an AI system. There are two ways an AI system can exceed its knowledge limits. One is if the query or operation is outside the domain of the AI system. As an example, in CF, successful file fragment classification becomes challenging for an AI algorithm when an image is embedded in a file that itself is not an image, e.g., an image embedded in a PDF file [70]. In this case, if the AI algorithm classifies the fragment as an image rather than a PDF file without any explanation, then the system using the AI algorithm will not be able to dismiss this as a misleading output. But if the AI algorithm knows its limitations and explains then the AI system will be able to dismiss this misleading output. The second is if the confidence of the output is too low. Revisiting the above example, classifying a fragment in a multi-class (several file types) is also a challenging problem [70]. In this case, if the AI algorithm classifies a fragment into more than one class and randomly chooses one class out of these as an output, then this will be a misleading output. An AI system with the knowledge of these limitations will not trust such output.
@@ 4.1.1 Classification of XAI Methods
We divide and classify the XAI methods by the scope of their explanation as shown in Figure 5. The scope of an explanation refers to its range and extent. It can be either local or global. Both local and global can be intrinsic and post-hoc. By definition, intrinsic methods are model-specific, and post-hoc methods are generally model-agnostic. This section presents some of the popular local and global methods. For a comprehensive coverage on this specific subject, readers are referred to [71, 72, 26].
@@
    Local Explanations – The scope of these explanations is a single input or a subset of inputs. The most common is the explanation for the output of a single input point. Some of the local explanation algorithms are LIME [44], SHAP [45, 46], Saliency Map [73, 74], and Counterfactual [75]. LIME explains each prediction of the original model by building a proxy self-interpretable model, such as linear models or decision trees. This new model is trained on a new dataset containing the interpretable representations of the original data. The original model can be explained by the weights given to features in the new model. For example, a model predicts that a patient has the flu. LIME in the explanation highlights the symptoms with relative weights, such as sneezing and headache are given more weight than no fatigue. A doctor can easily decide with the help of these explanations. SHAP also builds and uses a simpler model that is the interpretable approximation of the original model. It provides feature importance (contribution of each feature to the prediction result) by computing the Shapley value for each feature based on game theory. Saliency Map, also known as sensitivity map or heat map, explains each prediction by computing an attribute value for each feature based on its degree of influence on the prediction result. A saliency map is a visual explanation of a model where contributions can be visualized as heat maps. Counterfactual explanations indicate what amount of change in the input can modify the output. The objective is to minimize the new input value and bring it as close to the original input as possible. The goal is to maximize this objective function by a local search. This function depends on the endpoint loss, i.e., how close its value is to the desired class and the distance between the origin and the new value. Counterfactual explanations are similar to adversarial examples, but they differ in purpose. Unlike adversarial examples, where the purpose is to evade a classifier, they facilitate understanding of a model by providing explanations.
@@
    Global Explanations – The scope of these explanations is the entire algorithm. Some of the global explanation algorithms are PDP (Partial Dependence Plot) [76], ICE (Individual Conditional Expectation) [77], GSA (Global Sensitivity Analysis) [78, 79], SP-LIME (Submodular Pick LIME) [44]. PDP explanations indicate the marginal change of the output when there is a change in the feature value. The PDP focuses on the overall average instance rather than a specific one. It reveals the relationship between a feature and the prediction results. ICE plots are the graphs with line charts that represent the functional relationship between a feature and the predicted response. All the other features are kept fixed, only changing the value of the feature under analysis. GSA method perturbs the input features through their range of values and quantifies its effects on the predictions of a given model. SP-LIME, a variant of LIME, chooses the most relevant local LIME explanations to summarize and provide a global explanation.
@@ 4.2 Last Ten Years of XAI
To make readers familiar with the study carried out in the paper, we present a critical and short review of the last ten years of work carried out in XAI. We exclude papers that explore the application of XAI in CF because we discuss in detail such applications in section 6. A detailed survey on this topic is out of the scope of this paper. For a detailed survey, interested readers are referred to [30, 80].
@@
Hara et al. [81] propose a post-processing method to enhance the interpretability of Additive Tree Models (ATMs) [82] by approximating them with simpler, more human-interpretable models. The challenge addressed is the high number of regions that ATMs typically create, hindering interpretability. Mishra et al. [83] extend the LIME technique to music content analysis (MCA) named SLIME, introducing three versions of explanations based on temporal, frequency, and time-frequency segmentation. SLIME provides insights into model behavior, revealing limitations in the decision tree model’s generalization despite high accuracy and demonstrating model-agnostic explanations for the neural network, aligning with model-dependent saliency maps in many cases. Fong et al. [84] propose a model-agnostic framework for learning explanations for any black box algorithm, with a focus on identifying the image regions most responsible for classifier decisions. The proposed method employs explicit and interpretable image perturbations, addressing the limitations of heuristic-based image saliency methods. The framework contributes to interpretable and testable explanations, revealing insights into neural network fragility and susceptibility to artifacts. Ribeiro et al. [85] propose a model-agnostic explanation system called Anchors, which provides high-precision rules in the form of if-then statements to elucidate the behavior of complex models. Anchors efficiently compute explanations for any black-box model with high probability, showcasing their flexibility across various domains and tasks. A user study demonstrates that Anchors enable users to predict model behavior on unseen instances with higher precision and less effort compared to existing model-agnostic explanation techniques. Guidotti et al. [86] introduce LORE (for LOcal Rule-based Explanations), a model-agnostic method for explaining black box decisions, focusing on providing interpretable and faithful explanations for specific instances. LORE employs a genetic algorithm to generate a synthetic neighborhood and then builds a local interpretable predictor in the form of a decision tree. The resulting explanation includes a decision rule and a set of counterfactual rules. LORE outperforms existing methods in terms of explanation quality and accuracy in mimicking the black box, addressing the ethical issue of a lack of transparency in decision systems.
@@
Joo et al. [87] investigate the application of Grad-CAM to Deep Reinforcement Learning (DRL) in the context of Atari games. By revealing the significance of the CNN layer, the research enhances our understanding of how AI agents make decisions while playing Atari games through the application of Grad-CAM to DRL. Gramegna et al. [88] propose an XAI model for deciphering customer decisions in the non-life insurance domain. The authors integrate XGBoost with Shapley Values, which enhances behavioral segmentation, showcasing the potential of explainable machine learning to refine our understanding of customer behavior in the insurance industry. Cavaliere et al. [89] address the challenge of poor explainability in complex machine learning models, particularly in medical applications. Their research focused on automatic Parkinson’s disease detection, utilizing Grammar Evolution (GE) with Automatically Defined Functions. Parsa et al. [90] apply the eXtreme Gradient Boosting (XGBoost) machine learning technique to enhance traffic safety through rapid accident detection. The study demonstrated XGBoost’s effectiveness with high accuracy and detection rates, emphasizing the significance of traffic-related features, especially speed variations, through SHAP analysis. Feichtner et al. [91] aims to identify inconsistencies between developer-defined application behavior and permission usage in Android applications. It utilizes a deep neural network trained on descriptive texts to predict permission scores, with LIME employed for explainability analysis. De et al. [92] propose a method to enhance the interpretability of decisions made by Deep Neural Networks (DNNs) by combining Cluster-TREPAN, a decision tree generation algorithm, with hidden-layer-clustering. The focus is on explaining predictions from neural networks, emphasizing the importance of interpretability and reliability. The proposed Cluster-TREPAN method outperforms LIME in performance, effectively capturing and explaining the information flow within DNNs, addressing a critical aspect for human acceptance of deep learning models.
@@
Szczepański et al. [93] focused on enhancing the explainability of BERT-based fake news detectors, addressing the crucial need for understanding model decisions in real-life scenarios. Utilizing surrogate-type explainability methods, specifically LIME and Anchors, the study demonstrates the feasibility of enhancing interpretability in linguistics-based fake news detection models. Gite1 et al. [94] utilize an advanced combination of efficient machine learning techniques, specifically Long Short Term Memory (LSTM) for stock price prediction. The paper introduces an XAI element with LIME for transparent and meaningful insights into the model’s decision-making process. Naeem et al. [95] address the critical issue of automated malware detection in IoT devices, leveraging deep learning techniques, and using Grad-CAM to understand and improve the overall performance. Sutton et al. [96] apply the Grad-CAM technique to enhance the diagnostic precision of ulcerative colitis, utilizing 8000 endoscopic images. While the choice of Grad-CAM provides visual insights into model learning, a more detailed discussion of the AI techniques used and the specific XAI methods employed would provide a comprehensive evaluation of the study’s contributions. Debjit et al. [97] work presents an advanced machine learning framework for early COVID-19 detection, leveraging the Harris Hawks Optimization algorithm to optimize ML classifiers’ hyperparameters and SHAP values for model interpretation. Utilizing SHAP, the research offers transparency by elucidating feature importance, contributing to critical clinical insights in the domain of healthcare monitoring for COVID-19. Nordin et al. [98] propose an explainable predictive model for suicide attempts, combining complex ensemble learning models, such as Random Forest and Gradient Boosting, with the SHAP method. This approach aims to enhance reliability in predicting suicide attempts while offering insights into critical factors for clinical decision-making. In the domain of mental health, the study highlights the superiority of Gradient Boosting with SHAP over Random Forest in terms of accuracy, underscoring the value of explainable machine learning techniques in healthcare.
Ahmed et al. [99] introduce a hybrid adaptive ensemble that merges several classifiers and incorporates XAI components to produce transparent, feature-based explanations. Their approach employs post-hoc feature-importance methods and local explanation techniques, allowing analysts to identify which network-traffic attributes most strongly influenced a given prediction. The study in [100] addresses a major barrier to clinical AI adoption: determining whether current visual XAI tools genuinely support radiologists in reliably understanding model decisions. Instead of proposing a new method, the authors conduct a human-centered empirical assessment of widely used techniques, such as Grad-CAM, saliency maps, and other attribution visualizations, applied to chest X-ray classification. Their findings show that XAI visualizations vary in interpretability across different pathologies. Tahir et al. [101] present a hybrid XAI framework that combines LIME and SHAP to deliver instance-level interpretability while preserving the global coherence offered by Shapley-based explanations. This work represents a meaningful advance toward practical interpretable AV systems through the use of hybrid XAI strategies. Vani et al. [102] propose a deep learning approach for personalized health monitoring, employing SHAP to provide patient-specific interpretability. This enhances model transparency and promotes clinically grounded trust. Alamro et al. [103] develop an intrusion-detection system based on dimensionality reduction and attention-driven deep learning, integrating XAI to improve threat detection and support decision-making. SHAP is used to deliver reliable interpretive insights within AI-enabled security environments. Ghosh [104] combines XAI with a deep learning model to examine large-scale financial datasets. By capturing temporal patterns in network activity, logs, and audit trails, the system identifies fraudulent financial transactions. LIME is employed to ensure transparency and evaluate the model’s interpretability.